# RiteWeight: Randomized Iterative Trajectory Reweighting for Steady-State Distributions Without Discretization Error

### v3

@@ 3.2.2 Nonequilibrium Analysis
Trajectory preparation: We harnessed the complete equilibrium-like 208 s MD trajectory to generate trajectories for nonequilibrium steady state analysis. No subsampling of the trajectory data was performed. Instead, we extracted all possible two-step segments based on a lag time of 100, 10, 1, or 0.2 ns (the shortest lag time in the original MD data).
Macrostate definitions: Following prior work [25], we used the fuzzy spectral clustering method PCCA++ with a lag time of 100 ns to probabilistically assign configurations to folded and unfolded states. Then, we defined the “intermediate” region as the 10% of configurations whose membership scores were closest to 50% folded / 50% unfolded. We assigned the remaining 90% of configurations to either the folded or unfolded region based on the membership scores. These macrostates were used for both MSM and RiteWeight analysis.
MSM Analysis: Optimal MSM hyperparameters were previously identified for the 208 s MD data set [20, 35]. We used the PyEMMA software package [34] and replicated the optimal parameter choices, including features based on minimal residue-residue distances, dimensionality reduction using tICA with 100 tICs at a lag time of 10ns (with tICs rescaled using commute distances), and clustering using -means with .
RiteWeight Analysis: In each RiteWeight iteration, we performed clustering in the intermediate region using random clusters based on the same tICA coordinates used for MSM analysis. We then constructed a transition matrix that included the intermediate clusters along with the source and sink. To enforce the non-equilibrium steady-state (NESS) condition, we applied a boundary condition during the construction of the transition matrix. Specifically, the transition probability from the sink to all other clusters was set to zero, except for the source-state cluster which was set to one. We used the same source and sink definitions for RiteWeight and MSM analysis.
@@ 4.1 Synthetic MD for Trp-cage
RiteWeight’s accuracy is equally high using 10 or 1,000 clusters, underscoring the algorithm’s robustness. However, RiteWeight’s convergence speed does depend on the number of clusters, with more clusters yielding faster convergence. See the KL-based convergence analysis in supplementary Fig. 7. Following this analysis, we used iterations for RiteWeight with 1000 clusters and iterations for RiteWeight with 10 clusters. In both cases, the trajectory weights reported here are averages over the final 1,000 iterations.
In contrast to the converged RiteWeight estimates, the single-shot reweighting that occurs during the first RiteWeight iteration exhibits discrepancies due to discretization error. Even with 1,000 clusters, single-shot reweighting does not accurately predict the true stationary probabilities of the Trp-cage SynMD microstates. The discrepancy here can be attributed to the lack of local equilibrium within the clusters: local equilibrium is necessary for the success of single-shot reweighting but not for RiteWeight.
4.2 Atomistic MD for Trp-cage: Equilibrium and nonequilibrium steady states
@@
Fig. 3 shows estimates of the equilibrium distribution calculated using RiteWeight and MSMs. The reference values come from a direct computation using the full 208 s Trp-Cage trajectory, and the standard errors are computed by dividing the MD trajectory into three 70 ns blocks and comparing estimates. RiteWeight matches closely with the reference, despite starting from a distribution far from equilibrium. Here, the RiteWeight distribution is an average over the final 10,000 iterations in a run of 100,000 iterations. See the KL-based convergence analysis in supplementary Fig. 8.
By contrast, MSM-based reweighting fails to recover the equilibrium distribution, even when using the same features as RiteWeight with as many as 50,000 clusters. The inability to recover the correct stationarity distribution is due to an unsatisfied assumption that weights are locally in equilibrium within each cluster. The poor performance persists even when the MSM lag time is set to 100 ns; see supplementary Fig. 9.
In the nonequilibrium setting, Fig. 4 shows that RiteWeight is similarly successful. In this case, the reference stationary measure comes from modifying the 208 s Trp-cage trajectory by removing configurations that were more recently in the folded than the unfolded state. Again RiteWeight corrects the imbalances in the starting distribution and delivers a faithful copy of the stationary distribution. Here, the RiteWeight distribution is an average over the final 10,000 iterations in a run of 400,000 iterations, based on the mean first passage time convergence analysis in supplementary Fig. 10.
@@ 4.3 Atomistic MD for Trp-cage: Mean first-passage time
Next we used RiteWeight to calculate the mean first-passage time (MFPT) for atomistic trp-Cage folding. We ran RiteWeight using source-sink boundary conditions with the unfolded region as the source and the folded region as the sink. We then evaluated the MFPT using the reciprocal flux relation [36, 37]:
where is the lag time for the trajectory segments and is the estimated probability of each trajectory segment . For comparison, we also performed an MSM analysis using the first-step relation:
@@
RiteWeight is able to reproduce the MFPT because it uses a self-consistent iteration that locally corrects the weights within each cluster. In contrast, an MSM implicitly uses uniform weights within each cluster without correction. The initial trajectories are distributed close to equilibrium, and the MSM does not account for the nonequilibrium distribution from the source-sink boundary conditions that is “tilted” with respect to equilibrium [38]. We conjecture that updating the weights is especially important at short lag times, since the trajectories contributing to the MFPT estimate are the peripheries of the clusters which is the region of greatest tilting.
@@ 4.4 Atomistic MD for Trp-cage: Net fluxes
A major goal of molecular simulation is understanding the “mechanisms” of a functional transition, here defined as the temporal sequences of molecules. We analyzed the mechanisms for atomistic Trp-cage folding using net fluxes:
where the cluster indices and are ordered so that the net flux is positive in the reference MD data. For comparison, we also performed an MSM analysis using the formula for the net fluxes involving the backward committor [25]:
@@
Fig. 6 shows that the net fluxes from RiteWeight closely match the MD reference for each lag time 0.2, 1, 10, or 100 ns. Thus, RiteWeight precisely describes the transition processes leading to folding, even processes with a lag time ns which is the shortest lag time available in the original MD data. In contrast, the net fluxes from MSMs exhibit large discrepancies, echoing a previously identified deficiency of the MSM approach [25]. At each lag time, the MSMs exhibit some net fluxes in the opposite direction of MD, implying a different temporal sequences of events. The MSM fluxes are most quantitatively accurate at the 100 ns lag time, which is 1-2 orders of magnitude longer than the transition-path time over which folding events typically take place [25]. Supplementary Figs. 11 and 12 report confidence intervals for the MD reference fluxes, revealing tht a range of MSM estimates fall outside the 95% confidence intervals at each lag time 10 ns.
@@ 5 Discussion and Conclusions
We have introduced RiteWeight, an algorithm for correcting distributions that deviate from equilibrium or nonequilibrium steady states. Although it is built on a framework of Markov state models (MSMs) that discretize phase space, RiteWeight iteratively achieves self-consistency among all possible discretizations. This new approach overcomes a key limitation of MSMs by correcting the distribution within each discrete region to conform to the steady state. It accurately evaluates the steady state without requiring the Markov property at the discrete level. When the internal distributions match the steady state, a wider array of observables — such as mean first passage times and net fluxes — can be calculated with high accuracy even at short lag times [25]. The robustness of RiteWeight is illustrated by the insensitivity to the number of clusters used during the iteration process (Fig. 2).
@@
Building on previous reports [25], we have probed the shortcomings of MSMs for identifying mechanisms of folding in the atomistic Trp-cage system (Fig. 6). Surprisingly, the MSMs built from MD reference data continue to show discrepancies from reference values even at lag times of 100 ns. At sufficiently long lag times, the system must become fully Markovian, but evidently at 100 ns, the net fluxes are still influenced by the initial distributions within the clusters. This sensitivity of the net fluxes contrasts with the relative robustness of MSM mean first passage time estimates that match the reference value at a lag time of 100 ns (Fig. 5).
There remain important tasks ahead to realize the full potential of RiteWeight. Most notably, the data analyzed here consists of large data sets with dense sampling in visited regions of phase space. In sparser data sets, the algorithm may need to rely on other regularization and smoothing strategies. For example, we can modify the RiteWeight update formula (2) to read
Here, is a learning rate hyperparameter within the interval . Choosing a smaller value may mitigate the impact of noise without changing the RiteWeight fixed point (Theorem 1 in the Appendix). Another valuable strategy that may improve the estimation of observables in key phase-space regions is adaptive sampling [40, 41]. Finally, because RiteWeight can be accurate at any lag time, the range of lag times could be explored to optimize numerical efficiency in equilibrium calculations.
We foresee diverse opportunities for applying RiteWeight in the future. To promote the longstanding goal of generating Boltzmann-weighted ensembles for proteins, RiteWeight could be applied to MD trajectories initiated with machine-learning approaches based on AlphaFold [15, 16, 17, 18] or trajectories initiated with nuclear magnetic resonance data [42, 43]. Further, RiteWeight could be used to iteratively calculate local mean first passage times and generate transition paths in combination with the weighted ensemble method [44, 45].
In conclusion, the new RiteWeight algorithm solves the problem of reweighting flawed initial data sets into desired steady state distributions. The iterative use of random clusters leads to quasi-continuous and accurate final distributions. Applications to equilibrium and nonequilibrium observables compare favorably to state-of-the-art methods, with a particular advantage in quantifying nonequilibrium net fluxes.
@@
The python package for running the RiteWeight algorithm is available at: https://github.com/ZuckermanLab/rite_weight. The data used in this study to evaluate RiteWeight’s performance is available upon request. The SynMD Trp-cage model used in this study is hosted at: https://github.com/jdrusso/SynD-Examples
@@ Acknowledgments
We thank Jeremy Copperman and John Russo for valuable discussions on the subject of reweighting trajectories. The authors are grateful for financial support from the NIH under Grant GM115805. We thank DE Shaw Research for the Trp-cage trajectory data.
@@
This section identifies the “fixed point” of the RiteWeight algorithm, which is defined as follows.
@@
A collection of trajectory weights with is a fixed point of RiteWeight if the weights stay the same during each RiteWeight iteration for any positive probability choice of partition.
In practice, we typically apply RiteWeight using a continuous state space and clusters defined by a random Voronoi tesselation. Here for simplicity we consider a finite state space with clusters defined by a random hyperplane tesselation. See [46] for background on random hyperplane tesselations, which are called “stable under iteration tesselations” in stochastic geometry.
@@
The RiteWeight clusters are generated through one or more iterations of the following procedure. Initially, there is a single cluster containing all the microstates. At each iteration, any cluster that contains at least two microstates is randomly split by a hyperplane into two new clusters as follows. First, we choose a uniformly random direction that is normal to the hyperplane
@@
Under Assumptions 1 and 2, the following main result shows that the RiteWeight fixed point is completely determined by the microstate transition matrix.
@@
Consider a finite state space (Assumption 1) and a collection of trajectories with weights satisfying . Define the associated microstate transition matrix with entries
and assume has a unique stationary measure. Here, means that segment begins in microstate , and means that segment ends in microstate . Then, under the random hyperplane model (Assumption 2), is a fixed point of RiteWeight if and only if the vector with entries
is the fixed point of , that is,
@@
First we check that equation (11) implies is a fixed point of RiteWeight. By assumption, the microstate transition matrix has a unique stationary measure, so it is irreducible and aperiodic. It follows that the cluster transition matrix is irreducible and aperiodic, so has a unique stationary measure also. Next, equation (11) implies
This confirms that is a fixed point of RiteWeight.
Next, we assume that is a fixed point of RiteWeight and check that equation (11) holds. Since is a fixed point, we must have for each cluster and consequently
By considering all positive probability random clusters , we arrive at the identity
@@
We observe that the first iteration of the random hyperplane tesselation generates a uniformly random direction that leads to distinct values for every distinct microstate with probability one. Hence, there must be a consistent ordering of the microstates, say , so that the event
@@
occurs with positive probability for each . With the microstates ordered in this way, the linear span of the columns of includes all the vectors , where is a standard basis vector. This shows has full column rank, completing the proof. ∎

### v4

@@ 3.2.2 Nonequilibrium analysis
Trajectory preparation. We harnessed the complete equilibrium-like 208 s MD trajectory to generate trajectories for nonequilibrium steady state analysis. No subsampling of the trajectory data was performed. Instead, we extracted all possible two-step segments based on a lag time of 100, 10, 1, or 0.2 ns (the shortest lag time in the original MD data).
MSM analysis and definition of “folded” and “unfolded” states. Previous work has identified optimal MSM hyperparameters for the 208 s MD data set [20, 2]. We replicated the optimal parameter choices with the PyEMMA software package [39], including features based on minimal residue-residue distances, dimensionality reduction using tICA with 100 tICs at a lag time of 10 ns, rescaled tICs based on commute distances, and clustering using -means with . Next, following the prior work [42], we applied the fuzzy spectral clustering method PCCA++ with a lag time of 100 ns to identify two extreme states representing “folded” and “unfolded” structures. PCCA++ also assigned folded and unfolded membership scores to each of the remaining MSM states.
RiteWeight analysis. In each RiteWeight iteration, we performed clustering using random clusters based on the same tICA coordinates used for MSM analysis. We only clustered configurations in the intermediate region, excluding the previously identified MSM folded and unfolded states. We then constructed a transition matrix that included the intermediate clusters along with the folded and unfolded states. To enforce the non-equilibrium source-sink condition, we set the transition probability from the source to the sink equal to one and set the transition probability from the source to all other clusters equal to zero.
Flux analysis. Followed prior work [42], we identified “macrostates” for flux analysis with both MSMs and RiteWeight. First we selected the intermediate MSM states whose PCCA++ membership scores were closest to 50% folded / 50% unfolded. Then we grouped the remaining MSM states into a folded and an unfolded macrostate based on their membership scores.
@@ 4.1 Synthetic MD for Trp-cage
RiteWeight’s accuracy is equally high using 10 or 1,000 clusters, underscoring the algorithm’s robustness. However, RiteWeight’s convergence speed does depend on the number of clusters, with more clusters yielding faster convergence. See the convergence analysis in Supplementary Fig. S1. Following this analysis, we used iterations for RiteWeight with 1000 clusters and iterations for RiteWeight with 10 clusters. In both cases, the trajectory weights reported here are averages over the final 1,000 iterations.
In contrast to the converged RiteWeight estimates, the single-shot reweighting that occurs during the first RiteWeight iteration exhibits discrepancies due to discretization error. Even with 1,000 clusters, single-shot reweighting does not accurately predict the true stationary probabilities of the Trp-cage SynMD microstates. The discrepancy here can be attributed to the lack of local equilibrium within the discrete clusters: local equilibrium is necessary for the success of single-shot reweighting but not for RiteWeight.
4.2 Atomistic MD for Trp-cage: Equilibrium and nonequilibrium
@@
Fig. 3 shows estimates of the equilibrium distribution along the slowest tIC (tIC1), calculated using RiteWeight and an MSM. Supplementary Fig. S3 similarly shows a 2D projection of these estimates onto the two slowest tICs. The reference values come from a direct computation using the full 208 s Trp-Cage trajectory. RiteWeight matches closely with the reference despite starting from a distribution far from equilibrium. Here, the RiteWeight distribution is an average over the final 10,000 iterations in a run of 100,000 iterations. See the convergence analysis in Supplementary Fig. S2.
By contrast, MSM-based reweighting fails to recover the equilibrium distribution, even when using the same features as RiteWeight with as many as 50,000 clusters. The inability to recover the correct stationarity distribution is due to an unsatisfied assumption that weights are locally in equilibrium within each MSM state. The poor performance persists even when the MSM lag time is set to 100 ns; see Supplementary Fig. S4.
The equilibrium stationary distribution is characterized by a detailed balance condition [50]. However, RiteWeight does not enforce detailed balance directly, as do the MSM tools used for data analysis in this study [39]. Nevertheless, Supplementary Fig. S5 shows that RiteWeight recovers detailed balance to a similar or greater extent than a MSM.
Fig. 4 shows that RiteWeight is also successful in the nonequilibrium setting. Here, the reference stationary measure comes from modifying the 208 s Trp-cage trajectory by removing configurations that were more recently in the folded than the unfolded state. Again RiteWeight corrects the imbalances in the starting distribution and delivers a faithful copy of the stationary distribution. Here, the RiteWeight distribution is an average over the final 10,000 iterations in a run of 400,000 iterations, based on the mean first passage time convergence analysis in Supplementary Fig. S6.
@@ 4.3 Atomistic MD for Trp-cage: Mean first-passage time
Next we used RiteWeight to calculate the mean first-passage time (MFPT) for atomistic trp-Cage folding. We ran RiteWeight using source-sink boundary conditions with the unfolded region as the source and the folded region as the sink. We then evaluated the MFPT using the reciprocal flux relation [18, 5]
where is the lag time and is the weight assigned to each trajectory segment . For comparison, we also performed an MSM analysis using the first-step relation
@@
RiteWeight is able to reproduce the MFPT because it uses a self-consistent iteration that locally corrects the weights within each cluster. The initial weights are close to equilibrium, but the nonequilibrium distribution from the source-sink boundary conditions is “tilted” with respect to equilibrium [44]. We conjecture that updating the weights is especially important at short lag times because the trajectories determining the MFPT estimate are the peripheries of the clusters where the tilting is greatest.
@@ 4.4 Atomistic MD for Trp-cage: Net fluxes
A major goal of molecular simulation is understanding the “mechanisms” of a functional transition, here defined as the temporal sequences of the molecule. We analyzed the mechanisms for atomistic Trp-cage folding using net fluxes,
where the macrostate indices and were ordered to produce a positive net flux in the reference MD data. We analyzed fluxes involving a folded macrostate, an unfolded macrostate, and five intermediate macrostates, as described in Sec. 3.2.2 (Flux analysis). Supplementary Fig. S7 shows the intermediate macrostates projected onto the two slowest tICA components, and it illustrates the sequences of transitions from the unfolded to the folded macrostate.
For comparison, we also performed an MSM analysis using the formula for the net fluxes involving the backward committor [42]:
@@
Fig. 6 shows that the net fluxes from RiteWeight closely match the MD reference for each lag time 0.2, 1, 10, or 100 ns. Thus, RiteWeight precisely describes the transition processes leading to folding, even processes with a lag time ns which is the shortest lag time available in the original MD data. In contrast, the net fluxes from MSMs exhibit large discrepancies, echoing a previously identified deficiency of the MSM approach [42]. At each lag time, the MSMs exhibit some net fluxes in the opposite direction of MD, implying a different temporal sequences of events. The MSM fluxes are most quantitatively accurate at the 100 ns lag time, which is 1-2 orders of magnitude longer than the transition-path time over which folding events typically take place [42]. Supplementary Figs. S8 and S9 report confidence intervals for the MD reference fluxes, revealing tht a range of MSM estimates fall outside the 95% confidence intervals at each lag time 10 ns.
@@ 5 Discussion and conclusions
We have introduced RiteWeight, an algorithm for correcting distributions that deviate from equilibrium or nonequilibrium steady states. Although it is built on a framework of Markov state models (MSMs) that discretize configuration space, RiteWeight iteratively achieves self-consistency among all possible discretizations. This new approach overcomes a key limitation of MSMs by correcting the distribution within each discrete region to conform to the steady state. It accurately evaluates the steady state without requiring the Markov property at the cluster level. When the internal distributions match the steady state, a wider array of observables — such as mean first passage times and net fluxes — can be calculated with high accuracy even at short lag times [42]. The robustness of RiteWeight is illustrated by the insensitivity to the number of clusters used during the iteration process (Fig. 2).
Although RiteWeight employs coarse random clustering, over multiple iterations it adjusts the relative weights of trajectories at a fine resolution in feature space. Nevertheless, extremely close trajectories are unlikely to be separated into different clusters, so their relative weights are unlikely to change. To ensure the correction of weights, the feature space should be defined to differentiate configurations that are separated by a significant energy barrier. For example, in the Trp-Cage system, we constructed a suitable feature space using time-lagged independent components (tICs) derived from the dynamical behavior of the observed trajectories.
@@
The main requirement of the current implementation is that trajectory dynamics should be unbiased, i.e., outcomes from a given starting point should reflect the distribution that would occur in “ordinary” simulation without any unphysical forces or imposed stopping criteria. Thus, for example, weighted ensemble (WE) data could be used but only for lag times up to the WE resampling time; for longer lag times, a statistical continuation analysis would be required [8]. However, it should also be possible to use Girsanov reweighting of biased trajectories, a strategy previously applied for MSM construction [13].
Building on previous work [42], we have probed the shortcomings of MSMs for identifying mechanisms of folding in the atomistic Trp-cage system (Fig. 6). Surprisingly, the MSMs built from MD reference data continue to show discrepancies from reference values even at lag times of 100 ns. At sufficiently long lag times, the system must become fully Markovian, but evidently at 100 ns, the net fluxes are still influenced by the initial distributions within the MSM states. The sensitivity of the net flux estimates contrasts with the relative robustness of MSM mean first passage time estimates that match the reference value at a lag time of 100 ns (Fig. 5).
There remain important tasks ahead to realize the full potential of RiteWeight. Most notably, the data analyzed here consists of large data sets with dense sampling in visited regions of configuration space. In sparser data sets, the algorithm may need to rely on other regularization and configuration-space smoothing [30]. In addition, we can modify the RiteWeight update formula (2) to read
Here, is a learning rate hyperparameter within the interval . Choosing a smaller value may mitigate the impact of noise without changing the RiteWeight fixed point (Theorem 1 in the appendix). Another valuable strategy that may improve the estimation of observables in key configuration-space regions is adaptive sampling [9, 19]. Finally, because RiteWeight can be accurate at any lag time, the range of lag times could be explored to optimize numerical efficiency in equilibrium calculations.
We foresee diverse opportunities for applying RiteWeight in the future. To promote the longstanding goal of generating Boltzmann-weighted ensembles for proteins, RiteWeight could be applied to MD trajectories initiated with machine-learning approaches based on AlphaFold [21, 31, 22, 37] or initiated with nuclear magnetic resonance data [1, 23]. Further, RiteWeight could be used to iteratively calculate local mean first passage times and generate transition paths in combination with the weighted ensemble method [3, 36].
In conclusion, the new RiteWeight algorithm largely solves the problem of reweighting flawed initial data sets into desired steady state distributions. The iterative use of random clusters leads to quasi-continuous and accurate final distributions. Applications to equilibrium and nonequilibrium observables compare favorably to state-of-the-art methods, with a particular advantage in quantifying nonequilibrium net fluxes.
@@
The python package for running the RiteWeight algorithm is available at https://github.com/ZuckermanLab/rite_weight. The data used in this study to evaluate RiteWeight’s performance is available at https://doi.org/10.5281/zenodo.18489245. The SynMD Trp-cage model used in this study is hosted at https://github.com/jdrusso/SynD-Examples
@@ Acknowledgments
We thank Jeremy Copperman and John Russo for valuable discussions on the subject of reweighting trajectories. We thank DE Shaw Research for the Trp-cage trajectory data. The authors are grateful for financial support from the NIH under Grant GM115805. The research reported in this publication used computational infrastructure supported by the Office of Research Infrastructure Programs, Office of the Director, of the National Institutes of Health under Award Number S10OD034224.
@@
This section identifies the fixed point of the RiteWeight algorithm, which is defined as follows.
@@
A sequence of trajectory weights with is a fixed point of RiteWeight if the weights stay the same during each RiteWeight iteration for each positive probability choice of partition.
In practice, we typically apply RiteWeight in a continuous state space with clusters defined by a random Voronoi tesselation. Here for simplicity our analysis considers a finite state space with clusters defined by a random hyperplane tesselation. See [29] for background on random hyperplane tesselations, which are called “stable under iteration tesselations” in stochastic geometry.
@@
The RiteWeight clusters are generated through one or more iterations of the following procedure. Initially, there is a single cluster containing all the microstates. At each iteration, any cluster that contains at least two microstates is randomly split by a hyperplane into two new clusters as follows. First, we choose a uniformly random direction
@@
Under Assumptions 1 and 2, the following main result shows that the RiteWeight fixed point is determined by the microstate transition matrix. The microstate transition matrix does not change with the RiteWeight iterations. Therefore, assuming RiteWeight converges, it can only converge to a single fixed point.
@@
Assume a finite state space (Assumption 1) and a random hyperplane model (Assumption 2). Consider a sequence of trajectories with weights satisfying , and define the microstate transition matrix with entries
Assume has a unique stationary measure. Then the sequence of weights is the fixed point of RiteWeight if and only if the vector with entries
is the stationary measure of , that is,
@@
First we check that equation (11) implies is the fixed point of RiteWeight. This happens if the weights appear as entries of the left leading eigenvector of the cluster transition matrix . By assumption, the microstate transition matrix has a unique stationary measure, so it has one closed communicating class. It follows that the cluster transition matrix has one closed communicating class, so has a unique stationary measure also. Next, equation (11) implies
This confirms that the left leading eigenvector has the correct entries, so is the fixed point of RiteWeight.
Next, we assume that is the fixed point of RiteWeight and check that equation (11) holds. Since is the fixed point, each positive-probability cluster must satisfy
By considering all positive probability clusters, we arrive at the identity
@@
We observe that the first iteration of the random hyperplane tesselation generates a uniformly random direction that leads to distinct values for every distinct microstate with probability one. Hence, there is an ordering of microstates, say , so that the event
@@
occurs with positive probability for each . With the microstates ordered in this way, the linear span of the columns of includes all the vectors , where is a standard basis vector. This shows has full column rank and equation (11) holds, completing the proof. ∎